OpenStudy (anonymous):

integral from sin^5xdx is? please explain how you got your answer so i can undestand the probelm. thanks.. :D

zepdrix (zepdrix):

\[\Large\bf\sf \int\limits \sin^5x\;dx\]We gotta solve this bad boy?

zepdrix (zepdrix):

Oh it's an `odd` exponent, so this will work out nicely!

zepdrix (zepdrix):

\[\Large\bf\sf \color{#DD4747 }{\sin^2x\quad=\quad 1-\cos^2x}\]We'll apply this identity to our problem.\[\large\bf\sf \int\limits\limits \sin^5x\;dx\quad=\quad \int\limits \sin x \sin^2x \sin^2x\;dx\quad=\quad \int\limits \sin x(\color{#DD4747 }{\sin^2x})^2\;dx\]

zepdrix (zepdrix):

\[\Large\bf\sf \int\limits \sin x(\color{#DD4747 }{1-\cos^2x})^2\;dx\]

zepdrix (zepdrix):

From here we can make a nice easy u-substitution. \[\Large\bf\sf u=\cos x\]

zepdrix (zepdrix):

Think you can solve it from there? Shouldn't be too bad after you get your du.

OpenStudy (anonymous):

\[\int\limits_{}^{}(1-u^{2})^{2}sinxdx\]

OpenStudy (anonymous):

\[\int\limits_{}^{}(1-U ^{2})^{2}dU\]

OpenStudy (anonymous):

isn't it?

OpenStudy (anonymous):

thx

zepdrix (zepdrix):

derivative of cosx is -sinx. I think our substitution should bring along a negative sign in our du.

OpenStudy (anonymous):

so.. \[-\int\limits_{}^{}(1-u ^{2})^{2}du\] \[-\int\limits_{}^{}1-2u ^{2}+u ^{4}du\]

OpenStudy (anonymous):

how about the sinx one?